OpenStudy (anonymous):

How do you find the antiderivative of 3^x ?

OpenStudy (anonymous):

The antiderivative is basically the same as the indefinite integral. Do you know the form of the solution to \[\int\limits_{}^{}3^{x}dx\]?

OpenStudy (anonymous):

I know how to find the derivative of that, which would be 3^x (ln3). But I'm not sure how to find the antiderivative of it.

OpenStudy (anonymous):

Okay. What exactly is an anti-derivative? The opposite of a derivative, right? So, we start by taking the derivative of the function which is f(x) = 3^x f'(x) = 3^x(ln3), so You integrate the value to get the antiderivative of the function which would be3^x/(log(3)) Alternatively, ( 3^x) / ln3 + C -------------------------------------- 3^x = e^ {ln (3^x) } = e ^ {x ln3} whose antiderivative is [e ^ {x ln3}] / ln3 +C or [3^x] / ln3 +C or just remember that the derivative of 3^x is (3^x)ln3

OpenStudy (anonymous):

Okay but why is it that when we integrate the derivative we divide it by ln3?

OpenStudy (anonymous):

I mean I understand from the explanation, but are you supposed to divide by that every time?

OpenStudy (anonymous):

I suppose that's the rule. You do the opposite of whatever you would to find its derivative.

OpenStudy (anonymous):

Okay thanks! :)